At what rate must php 60,000 be invested to earn php 3,600 in 120 days.
Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

Depends on the compounding period.

Answer by Edwin McCravy(20055)

Obviously you meant SIMPLE interest, not compound interest.

I = P∙R∙T

I = 3600
P = 60000
T = 120/360 = 1/3
R = ?

3600 = 60000∙R∙(1/3)

Multiply 60000 by 1/3, get 20000

3600 = 20000∙R

Divide both sides by 3600

3600/20000 = R

0.18 = R

Answer 18% simple interest.

Edwin